A company has an EPS of $2.25, a book value per share of $23.85, and a market/book ratio of 3.0×. What is its P/E ratio? Do not round intermediate calculations. Round your answer to two decimal places.

×
Answer #1

Market/book ratio=Market value/Book value

Hence market value per share=(23.85*3)=$71.55

P/E ratio=market price/EPS

=71.55/2.25

=31.8